OpenStudy (cedric):

Consider the functions f(x) = x^2 − 13 and g(x) = x + 5. What is the value of f[g(−4)]? −12 3 −13 8

OpenStudy (anonymous):

g(-4) = -4+5 = 1 f(1) = 1^2-13= -12 The first one

OpenStudy (cedric):

How did you know what number to put in the variable?

OpenStudy (anonymous):

because you are asked for \(f(g(-4))\) so first you must find \(g(-4)\) and then you find \(f\) of the result from the first calculation

OpenStudy (precal):

always work from the inside to the outside